Further maths question
1.Express 3x-1/(x+1)(x-3) in a partial fraction
2. Simplify 2/3+5
3.Solve the quadratic equation using factorization method
a. 4x²+8x+1=0
b.3x²+4x-3=0
Maths question
1.Using a ruler and a pair of compass only construct a triangle PQR such that /QR/=6.5cm, /PQ/=3.2cm. a.construct the locus L1 of point equidistance from B and Q. Join RS and PS where S is point of intersection of L1 and L2. Measure /QS/
1
Expert's answer
2020-04-28T14:51:35-0400

1.

"\frac{3x-1}{(x+1)(x-3)}=\frac{A}{x+1}+\frac{B}{x-3}"

"\frac{A}{x+1}+\frac{B}{x-3}=\frac{A(x-3)+B(x+1)}{(x+1)(x-3)}=\frac{3x-1}{(x+1)(x-3)}"


"A(x-3)+B(x+1)=3x-1"

"Ax-3A+Bx+B=3x-1"

"(A+B)x-(3A-B)=3x-1"


"\begin{alignedat}{2}
 A+B = 3 \\
 3A-B = 1
\end{alignedat}" "\begin{alignedat}{2}
 A = 1 \\
 B = 2
\end{alignedat}"

Answer

"\frac{1}{x+1}+\frac{2}{x-3}"


2. "\frac{2}{3}+5=5\frac{2}{3}=\frac{17}{3}"


3. a. "4x^{2}+8x+1=0"

"x^{2}+2x+0.25=0"

"x^{2}+2x+1-1+0.25=0"

"(x+1)^{2}-0.75=0"

"(x+1)^{2}=0.75"

"x_{1}+1=\sqrt{0.75}" "x_{2}+1=-\sqrt{0.75}"

"x_{1}=\sqrt{0,75}-1" "x_{2}=-\sqrt{0,75}-1"

"x_{1}=\frac{\sqrt{3}}{2}-1" "x_{2}=-\frac{\sqrt{3}}{2}-1"


b. "3x^{2}+4x-3=0"

"x^{2}+\frac{4}{3}x-1=0"

"x^{2}+\frac{4}{3}x-1+\frac{4}{9}-\frac{4}{9}=0"

"(x+\frac{2}{3})^{2}-\frac{13}{9}=0"

"(x+\frac{2}{3})^{2}=\frac{13}{9}"

"x_{1}+\frac{2}{3}=\sqrt\frac{13}{9}" "x_{2}+\frac{2}{3}=-\sqrt\frac{13}{9}"

"x_{1}=\frac{\sqrt{13}-2}{3}" "x_{2}=-\frac{\sqrt{13}+2}{3}"